OpenStudy (anonymous):

Sketch the solution to the initial value problem and determine its maximum value. dy/dt=2y-2yt y(0)=3

OpenStudy (anonymous):

Arrange the given equation. (dy/2y) = 1-t now integrate both side . \[\int\limits_{y0}^{y}dy/2y =\int\limits_{0}^{t}1-t\] (1/2)* log(y/y0) = t-(t^2/2) y0 is nothing but value of y at initial point of time at t=0 so y0=3 ; (from initial boundary condition) log(y/3) = 2*{t-(t^2/2)} y = 3*e^2*{t-(t^2/2)}  (1) y will max. when 2*{t-(t^2/2)} will be max. because base 'e' is >1. so next task is to maximize t-(t^2/2) t-(t^2/2) = t*(1-(t/2)) max at t=1, that is 0.5 put in eqn.(1) Ymax= 3*e^2